# DIFFender: Diffusion-Based Adversarial Defense against Patch Attacks

### v3

@@ 3.3 Prompt Tuning
Following the aforementioned pipeline, leveraging visual-language pre-training, DIFFender is capable of efficiently performing zero-shot patch localization. While it is accurate in locating adversarial regions in most cases, subtle discrepancies in the segmented masks may occur in certain situations. Given that visual-language pretraining takes advantage of large-capacity text encoders to explore a vast semantic space [43], to facilitate the effective adaptation of learned representations into adversarial defense tasks, we introduce the algorithm of prompt tuning.
Learnable prompts. First, we replace the textual vocabulary with the learnable continuous vectors. Thus, $prompt_{L}$ and $prompt_{R}$ can be transformed into vectors as follows:
@@
where each $[V_{S}]_{i}$ or $[V_{R}]_{i}$ ($i\in\{1,\ldots,n\}$) is a vector of the same dimensionality as word embeddings. $n$ is a hyperparameter that specifies the number of context tokens, we set $n=16$ by default. The text content used to initialize $prompt_{L}$ and $prompt_{R}$ can be manually provided or randomly initialized.
@@
First, to accurately identify the adversarial regions for the localization module, we employ cross-entropy loss by comparing estimated mask $\hat{\mathbf{M}}$ with ground-truth mask $\mathbf{M}$.
@@
Lastly, to verify that the adversarial effects have been eliminated, we draw inspiration from [19] and [41] to make the high-level feature representations of the downstream classifiers between the restored image $\mathbf{x}_{r}$ and the clean image $\mathbf{x}$ close to each other. Specifically, we compute the $\ell_{2}$ distance between their feature representations at each layer weighted by a layer-wise hyperparameter as
@@
By summing up the aforementioned three losses, we have the overall loss $L_{PT}$ for prompt tuning as
@@
We perform gradient descent to minimize $L_{CE}$ w.r.t. $prompt_{L}$ and $prompt_{R}$ for prompt tuning. The design of continuous representations also enables thorough exploration in the embedding space.
Few-shot learning. During prompt-tuning, we utilize a minimal number of images for few-shot tuning. Specifically, DIFFender is tuned on a limited set of attacked images (8-shot in the experiments) from a single scenario and a specific attack method, but can learn optimal prompts that generalize well to other scenarios and attacks, which makes the tuning module effective and straightforward.
@@ 4 Experiments
Datasets and network architectures. We consider ImageNet [5] for evaluation and compare with eight state-of-the-art defense methods: Image smoothing-based defenses, including JPEG [8] and Spatial Smoothing [38]; Image completion-based defenses, such as DW [11], LGS [23] and SAC [20]; Feature-level suppression defense FNC [40] and Jedi [31], a defense based on entropy and diffusion-based defense DiffPure [24]. For classifiers, we consider two advanced classifiers trained on ImageNet: CNN-based Inception-V3 [30] and Transformer-based Swin-S [30].
Adversarial attacks. We employ AdvP [2] and LaVAN [17], which randomly select patch positions and generate perturbations. We also employ GDPA [18], which optimizes the patch’s position and content to execute attacks, and RHDE [33], which utilizes realistic stickers and searches for their optimal positions to launch adversarial attacks. To adaptively attack the baseline based on preprocessing, we approximate gradients using the BPDA [1], which implies that the defense methods are white-box against the attack methods and set the number of iterations for the attacks to 100. For adapting the attack on DIFFender, we use an additional Straight-Through Estimator (STE) [39] during backpropagation through thresholding operations. Additionally, due to the randomness introduced by the diffusion processes, we use the Expectation over Transformation (EOT) + BPDA attack [1].
Evaluation metrics. We evaluate the performance of defenses under standard accuracy and robust accuracy. Due to the computational cost of adaptive attacks, unless otherwise specified, we assess the robust accuracy on a fixed subset of 512 sampled images from the test set. To facilitate the observation of changes, we ensure that the selected subset consists of images correctly classified.
@@ 4.2 Evaluation on ImageNet
Table 1: Clean and robust accuracy against patch attacks on ImageNet by Inception-V3 and Swin-S.
Inception-V3 Swin-S
@@
Figure 6: Visualization with examples from ImageNet. The restored images of DIFFender no longer show any traces of the patch, and the restored details are remarkable (e.g., the recovery of tree branches in the second column of images).
(1) DIFFender outperforms in defense effectiveness. Under adaptive attacks utilizing gradients, such as the BPDA+AdvP and BPDA+LaVAN, DIFFender exhibits superior performance, even only involving an 8-shot process. Other attacks like GDPA may not achieve strong attack effectiveness, but DIFFender still attains the highest robust accuracy. This can be attributed to that DIFFender is built upon the unified diffusion framework. Intuitively, the diffusion model can effectively remove adversarial areas while ensuring a high-quality and diverse generation that closely follows the underlying distribution of clean data. Additionally, the inherent stochasticity in the diffusion model allows for robust stochastic defense mechanisms [13], which makes it a well-suited “defender” for adversarial attacks.
@@
(3) In the experiments, DIFFender only undergoes 8-shot prompt-tuning specifically for the AdvP method. This demonstrates the generalization capability of DIFFender to handle unseen attack methods. For Jedi, it has strong robustness against several attack methods, such as AdvP, but its robust accuracy has significantly decreased under other methods, like LaVAN. This might be because the autoencoder used by Jedi is trained under a specific style and cannot generalize well.
(4) When defending against global perturbations with $\ell_{p}$-norm constraints, DiffPure achieves excellent results. However, it performs poorly when facing patch attacks. Specifically, in Tab. 1, when tested against AdvP and LaVAN, the inception-v3 model purified by DiffPure only maintains robust accuracy rates of 10.5% and 15.2%, respectively. This is also consistent with our observations in Section 3.1
@@ 4.3 Ablation studies and additional results
Table 3: Ablation study for different modules in DIFFender. DIFFender (NR) denotes “No Restoration”.
@@
Table 4: Ablation study for different prompt forms. “EP” and “MP” represent “Empty Prompt” and “Manual Prompt”.
@@
Table 5: DIFFender transferability performance on ResNet50 and ViT-B-16 for ImageNet.
ResNet-50 ViT-base
@@
Impact of Prompt Tuning. In Tab. 4 we compared the complete DIFFender with the “Empty prompt” and “Manual prompt” versions of DIFFender. For DIFFender with manual prompts, we set $prompt_{L}$ = “adversarial” and $prompt_{R}$ = “clean”. The prompt-tuned DIFFender shows a significant improvement in robust accuracy compared to the other two zero-shot DIFFenders. This improvement, despite exposure to only a few attacked images, underscores the effectiveness of prompt tuning.
Cross-model transferability. Specifically, we conduct 8-shot prompt-tuning on Inception-V3 and Swin-S. We then test the transferability of learned prompts on CNN-based ResNet50 [12] and transformer-based ViT-B-16 [7]. The results are presented in Tab. 5. DIFFender keeps high robust accuracy when applied to new classifiers. This demonstrates the good generalization capabilities of DIFFender.
4.4 Extension in Face Recognition.
Experimental settings. Facial expressions in human faces introduce a rich diversity, together with external factors such as lighting conditions and viewing angles, making face recognition a challenging task. We conducted experiments on the LFW [15], and employed two adversarial patch attacks: RHDE [33] and GDPA [18] mentioned above.
Table 6: Clean and robust accuracy against patch attacks on LFW.
@@
Figure 7: DIFFender defense demonstrations against meaningless and meaningful patch attacks in the physical world. Noticing the mask edges may extend slightly beyond the patch region, aids in restoring the patch and helps maintain consistency in the restored image.
Figure 8: Visualization with examples from LFW attacked by RHDE, localized and restored by DIFFender.
Experimental results. The results on the LFW dataset are presented in Tab. 6. DIFFender achieves the highest robust accuracy under both the GDPA and RHDE while ensuring clean accuracy. It is worth noting that DIFFender is not re-tuned specifically for facial recognition. This further demonstrates the generalizability of DIFFender across different scenarios and attack methods. In contrast, JPEG, SS, and the FNC method obtained low robust accuracies, even below the clean accuracy. This is because in the specific context of facial recognition, the classifier focuses more on crucial local features, and preprocessing the entire image can disrupt these important features. Fig. 8 illustrates the results of DIFFender against face attacks. It can be observed that DIFFender accurately identifies the location of the patch and achieves excellent restoration.
@@
We additionally conduct further experiments in the physical world, where select 10 common object categories from ImageNet and perform two types of patch attacks (meaningful and meaningless) [34]. Our approach involves generating digital-world attack results first, then placing stickers on real-world objects in the same positions. We test DIFFender under various conditions, including different angles (rotations) and distances. Qualitative results are depicted in Fig. 7, and additional results can be found in Appendix A.5, while quantitative results are presented in Tab. 7, where each configuration is based on 256 frames successfully classified images from the 10 objects selected. Based on the results, we see that DIFFender manifests substantial defensive capabilities across various physical alterations, maintaining its efficacy in real-world scenarios.
Table 7: Quantitative accuracy of meaningless physical attacks on the Inception-v3 at different angles and distances, along with the quantized results after defense by DIFFender.
yaw ±15° yaw ±30° pitch ±15° distance Undefended 28.9 34.8 41.8 36.7 35.9 DIFFender 80.9 76.6 77.7 75.4 73.8
@@ 5 Discussion and Conclusion
We propose DIFFender, a novel defense method that leverages a pre-trained unified diffusion model to perform both localization and restoration of patch attacks. Additionally, we design a few-shot prompt-tuning algorithm to facilitate simple and efficient tuning. To show the robust performance of our method, we conduct experiments on image classification, further validate our approach on face recognition and finally extend to the physical world. Our research findings demonstrate that DIFFender exhibits superior robustness even under adaptive attacks and extends the generalization capability of pre-trained large models to various scenarios, diverse classifiers, and multiple attack methods, requiring only a few-shot prompt-tuning. We prove that DIFFender significantly reduces the success rate of patch attacks while producing realistic restored images.

### v4

@@ 3.3 Prompt Tuning
Following the aforementioned pipeline, leveraging vision-language pre-training, DIFFender is capable of efficiently performing zero-shot localization and restoration. While it is accurate in locating and restoring in most cases, subtle discrepancies may occur in certain challenging situations. Given that vision-language pre-training takes advantage of large-capacity text encoders to explore a vast semantic space [52], to facilitate the effective adaptation of learned representations into adversarial defense tasks, we introduce the algorithm of prompt tuning.
Learnable prompts. First, the textual vocabulary is replaced by learnable vectors. Thus, $prompt_{L}$ and $prompt_{R}$ can be transformed into vectors as follows:
@@
where each $[V_{L}]_{i}$ or $[V_{R}]_{i}$ ($i\in\{1,\ldots,n\}$) is a vector of the same dimensionality as word embeddings. $n$ is a hyperparameter that specifies the number of context tokens, we set $n=16$ by default. The text content used to initialize $prompt_{L}$ and $prompt_{R}$ can be manually provided or randomly initialized.
@@
First, to accurately identify the adversarial regions, we employ cross-entropy loss by comparing estimated mask $\hat{\mathbf{M}}$ with ground-truth mask $\mathbf{M}$.
@@
Lastly, to verify that the adversarial effects have been eliminated, we draw inspiration from [23] and [49] to make the high-level feature representations of the downstream classifiers between the restored image $\mathbf{x}_{r}$ and the clean image $\mathbf{x}$ close to each other. Specifically, we compute the $\ell_{2}$ distance between their feature representations weighted by a layer-wise hyperparameter as
@@
By summing three losses, the overall prompt tuning loss $L_{PT}$ is:
@@
We perform gradient descent to minimize $L_{PT}$ w.r.t. $prompt_{L}$ and $prompt_{R}$ for prompt tuning. The design of continuous representations also enables thorough exploration in the embedding space.
Few-shot learning. During prompt tuning, we utilize a limited number of images for few-shot tuning. Specifically, DIFFender is tuned on a limited set of attacked images (8-shot in the experiments) from a single attack, but can learn optimal prompts that generalize well to other scenarios and attacks, which makes the tuning module effective and straightforward.
@@ 4 Experiments
Datasets and Baselines. We consider ImageNet [6] for evaluation against eight state-of-the-art defense methods: Image smoothing-based defenses, including JPEG [11] and Spatial Smoothing [46], image completion-based defenses such as DW [14], LGS [29] and SAC [24], feature-level suppression defense FNC [48], alongside Jedi [37], a defense based on entropy. Additionally, we assess the diffusion purification defense, DiffPure [30]. For classifiers, we consider two advanced classifiers: CNN-based Inception-v3 [35] and Transformer-based Swin-S [35].
Adversarial attacks. We employ AdvP [2] and LaVAN [20], which randomly select positions and optimize patches. GDPA [22], which optimizes the patch’s position and content to execute attacks, and natural-looking attack RHDE [39], which utilizes realistic stickers and searches for their optimal positions to launch attacks. To implement adaptive attacks, we approximate gradients using the BPDA [1] to conduct BPDA+AdvP and BPDA+LaVAN attacks, which implies that the defense methods are white-box against the attacks. The number of iterations for the attacks is set to 100 with patch size 5% of the input image. For adapting the attack on DIFFender, we use an additional Straight-Through Estimator (STE) [47] during backpropagation through thresholding operations.
Evaluation metrics. We evaluate the performance of defenses under standard accuracy and robust accuracy. Due to the computational cost of adaptive attacks, unless otherwise specified, we assess the robust accuracy on a subset of 512 sampled images from the test set. To facilitate the observation, we ensure that the selected subset consists of images correctly classified.
@@ 4.2 Evaluation on ImageNet
Table 1: Accuracy (%) against attacks on ImageNet by Inception-v3 and Swin-S.
Models Inception-v3 Swin-S
Adaptive Non-adaptive Adaptive Non-adaptive
@@
Figure 6: Visualization with examples from ImageNet. The restored images of DIFFender exhibit no residual signs of the adversarial patch, and the restored details are remarkable (e.g., the recovery of tree branches in the second column of images).
@@
(1) DIFFender outperforms in defense effectiveness. Under adaptive attacks utilizing gradients, such as the BPDA+AdvP and BPDA+LaVAN, DIFFender exhibits exceptional performance, even only involving an 8-shot process. Other attacks like GDPA may not achieve strong attack effectiveness, but DIFFender still attains the highest robust accuracy. This can be attributed to that DIFFender is built upon the unified diffusion framework. Empowered by the AAP phenomenon, the diffusion model can effectively locate and remove adversarial areas while ensuring a high-quality and diverse restoration that closely follows the underlying distribution of clean data. Additionally, the inherent stochasticity in the diffusion model allows for robust stochastic defense mechanisms [16], which makes it a well-suited "defender" for adaptive attacks.
@@
(3) DIFFender demonstrates generalizability to unseen attacks. In the experiments, DIFFender only undergoes 8-shot prompt tuning specifically for the AdvP method yet achieves promising results under other attacks as well. For Jedi, it has strong robustness against several attack methods, such as AdvP, but its robust accuracy has significantly decreased under other methods, like LaVAN. This might be because the autoencoder used by Jedi is trained under a specific style and cannot generalize well.
(4) For the naturalistic attack RHDE, it poses a lesser threat to classifiers compared to adaptive meaningless attacks. However, it introduces a more significant challenge to defense methods, likely due to its utilization of irregular, more naturally-appearing patches. Nevertheless, DIFFender still achieves the best defense results without prior exposure to RHDE patches. Moreover, DIFFender is adaptable; with the prompt tuning module, a few-shot tuning can be employed to enhance performance specifically against naturalistic patch attacks.
(5) When defending against global perturbations with $\ell_{p}$-norm constraints, DiffPure achieves excellent results. However, it performs poorly when facing patch attacks. Specifically, in Tab. 1, when tested against AdvP and LaVAN, the Inception-v3 model purified by DiffPure maintains robust accuracy rates of 10.5% and 15.2%, respectively. This aligns with our observations in Sec. 3.1.
@@ 4.3 Ablation studies and additional results
Impact of patch size. We conducted experiments under patch attacks of varying sizes, using patches generated by AdvP ranging from 0.5% to 15% in size, and compared them with the state-of-the-art SAC and Jedi. As shown in Tab. 3, DIFFender exhibits better generalization to patches of various sizes, benefitting from vision-language pre-training, whereas Jedi and SAC are more sensitive to patch size. Notably, DIFFender was only prompt-tuned for patches of 5.0% size.
@@
Table 3: Accuracy against attacks of varying patch sizes by Inception-v3.
Size 0.5% 1.0% 5.0% 10.0% 15.0%
Undefended 64.3 50.8 0.0 0.0 0.0
SAC 81.8 83.8 84.2 60.9 34.8
Jedi 61.7 56.4 67.6 42.2 33.8
DIFFender 86.1 87.3 88.3 70.5 56.6
Table 4: Ablation study for restoration modules in DIFFender. "NR" denotes "No Restoration Process".
@@
Table 5: Ablation study for different prompt forms. "EP" and "MP" represent "Empty Prompt" and "Manual Prompt".
@@
Table 6: Transferability of DIFFender on ResNet50 and ViT-B-16 for ImageNet.
ResNet-50 ViT-B-16
@@
Impact of Prompt Tuning. In Tab. 5, DIFFender with prompt-tuning is compared with the "Empty prompt" and "Manual prompt" versions of DIFFender. For DIFFender with manual prompts, we set $prompt_{L}$ = "adversarial" and $prompt_{R}$ = "clean". The prompt-tuned DIFFender shows a notable improvement in robust accuracy compared to the other two zero-shot versions, despite exposure to only a few attacked images, underscoring the effectiveness of prompt-tuning.
Cross-model transferability. We apply only 8-shot prompt-tuning exclusively on Inception-v3. Subsequently, the transferability of DIFFender is tested on diverse classifiers, including the CNN-based ResNet50 [15] and Transformer-based ViT-B-16 [10]. The results are detailed in Tab. 6, underscore DIFFender’s ability to maintain robust accuracy across novel classifiers, demonstrating its potent generalization capacity.
Figure 7: Visualization with examples from LFW attacked naturally by RHDE, localized and restored by DIFFender.
Table 7: Accuracy against patch attacks on LFW by FaceNet.
@@
4.4 Extension in Face Recognition.
Experimental settings. Facial expressions in human faces introduce a rich diversity, together with external factors such as lighting conditions and viewing angles, making face recognition a challenging task. We conducted experiments on the LFW [18], and employed two patch attacks: RHDE [39] and GDPA [22].
Experimental results. The results on the LFW dataset are presented in Tab. 4.3. DIFFender achieves the highest robust accuracy under both the GDPA and RHDE while ensuring clean accuracy. It is important to note that DIFFender is not tuned specifically for facial recognition. This further demonstrates the generalizability of DIFFender across different scenarios and attack methods. In contrast, JPEG, SS, and the FNC method obtained low robust accuracies. This is because in the specific context of facial recognition, the classifier focuses more on crucial local features, and preprocessing the entire image can disrupt these important features. Fig. 7 illustrates the results of DIFFender against face attacks. It can be observed that DIFFender accurately identifies the location of the natural patch and achieves excellent restoration.
@@
We additionally conduct further experiments in the physical world, where we select 10 common object categories from ImageNet and perform two types of patch attacks (natural and meaningless) [40]. Our approach involves generating digital-world attack results first, then placing stickers on real-world objects in the same positions. We test DIFFender under various conditions, including different angles (rotations) and distances. Qualitative results are depicted in Fig. 8, while quantitative results are presented in Tab. 8, where each configuration is based on 256 frames successfully classified images from the 10 objects selected. Based on the results, we see that DIFFender demonstrates substantial defensive capabilities across various physical alterations, maintaining its efficacy in real-world scenarios.
Figure 8: Physical world defense demonstrations of DIFFender against meaningless and natural patch attacks. The mask edges may extend slightly beyond the patch region, aid in restoring the patch, and help maintain consistency in the restored image.
Table 8: Quantitative result of meaningless physical attacks on the Inception-v3 at different angles and distances.
yaw ±15° yaw ±30° pitch ±15° distance Undefended 28.9 34.8 41.8 36.7 35.9 Jedi 61.7 57.8 66.4 63.3 62.1 DIFFender 80.9 76.6 77.7 75.4 73.8
@@ 5 Discussion and Conclusion
We propose DIFFender, a novel defense framework harnessing a pre-trained unified diffusion model for dual roles in the localization and restoration of patch attacks, empowered by the discovery of the Adversarial Anomaly Perception (AAP) phenomenon. Additionally, we design a few-shot prompt-tuning algorithm to facilitate simple and efficient tuning, thus eliminating the need for extensive retraining. To validate the robust performance of DIFFender, we conduct experiments on image classification, face recognition, and further the physical world scenarios. Our findings demonstrate that DIFFender exhibits exceptional robustness even under adaptive attacks and extends the generalization capability of pre-trained large models to various scenarios, diverse classifiers, and multiple attack methods, requiring only a few-shot prompt-tuning. We prove that DIFFender effectively reduces the success rate of patch attacks while producing realistic restored images, promising a wide spectrum of diffusion model applications and inspiring future explorations in the domain.
There are several avenues for further exploration. While we have designed acceleration techniques to expedite diffusion-based methods, further acceleration can be achieved by adopting advanced acceleration sampling methods. Another potential direction for exploration is to extend the DIFFender framework to other modalities of adversarial attack defenses, such as video data.
Acknowledgement
This work was supported by the Project of the National Natural Science Foundation of China (No. 62076018, 92370124, 62276149, 92248303), the Fundamental Research Funds for the Central Universities, and Tsinghua-Alibaba Joint Research Program. Y. Dong was also supported by the China National Postdoctoral Program for Innovative Talents and CCF-BaiChuan-Ebtech Foundation Model Fund.